OpenStudy (anonymous):

The minimum number in a set of data is -2. If the range of the data is 10, what is the maximum number in the set?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

there are no numbers in the question :/

OpenStudy (anonymous):

except the -2 and 10

OpenStudy (anonymous):

Do you understand the question?

OpenStudy (ghazi):

12

OpenStudy (anonymous):

how'd u get that?

OpenStudy (ghazi):

range is 10 and least value is -2, so total number of data is 10 and 2 more sorry i guess its 13, including 0

OpenStudy (anonymous):

@nincompoop verify please?

OpenStudy (nincompoop):

range = maximum value - minimum value 10 = x - (-2) 10 = x + 2 10 - 2 = x + 2 - 2 8 = x So, our maximum value is 8.

OpenStudy (anonymous):

Oh, thanks!

OpenStudy (anonymous):

So it was actually range they were asking for?

OpenStudy (nincompoop):

make sure you look at your notes on maximum value, minimum value and range. I think this is some sort of getting you into the language of statistics.

OpenStudy (nincompoop):

The range and minimum value were given. You are being asked to solve for the maximum value. The "formula" is Range = maximum value - minimum value This problem wants you to look for the maximum value. We just did a little bit of "algebraic" manipulation to get it.

OpenStudy (anonymous):

Hmmm. Well, thank you!

OpenStudy (nincompoop):

Do you grasp what I did and why I did it that way?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

I didn't know there was a formula for it.

OpenStudy (anonymous):

BRB.

OpenStudy (anonymous):

Will go ahead and post my next question, then I will come back

OpenStudy (nincompoop):

Let me elaborate. A data set 4, 8, 12, 16, 18, 22 has maximum value of 22 and minimum value of 4. So if we were to look for the range we use the formula: range = maximum value - minimum value range = 22 - 4 range = 18 Let us switch things around and say that we are looking for the maximum value. range = maximum value - minimum value 18 = x - 4 (I use "x" but you can use any placeholder) we add 4 on both sides of the equation so x is left by itself. 18 + 4 = x - 4 + 4 22 = x Therefore, our maximum value is 22.

OpenStudy (nincompoop):

I am going to leave now. my lab has ended.

OpenStudy (nincompoop):

I will try again later when I get home

OpenStudy (anonymous):

Aw. Okay :(

OpenStudy (anonymous):

Will you come back later?

OpenStudy (nincompoop):

yeah because you and I have to do Spanish

OpenStudy (anonymous):

Great! :)

OpenStudy (anonymous):

Indeed.

OpenStudy (anonymous):

*hums*

OpenStudy (anonymous):

Oh, and just because of Spanish? LOL.

OpenStudy (anonymous):

Ok then. Bye.